# Evaluating GPT-4 at Grading Handwritten Solutions in Math Exams

### v1

@@ 1 Introduction
Automated scoring is a key challenge to enable the deployment of open-ended questions at scale. Researchers have studied the problems of automated essay scoring [1] and automated short-answer grading [3] extensively, often using AI. Fine-tuning language models, such as BERT, has been effective on these tasks [6], while recent works have explored prompting large language models for automated scoring [5]. One important setting in automated scoring is student handwriting: in many practice and assessment settings, students write down their solutions to problems on paper, which are then converted to images, common in science, technology, engineering, and math (STEM) fields [2]. This automated scoring task is challenging: compared to scoring textual essays and short answers, images contain rich semantic, visual, and even spatial information on student thought processes, which require significant textual, mathematical, and visual reasoning capabilities from AI. Recent advances in multi-modal foundation models, especially vision-language models, have significantly advanced the textual and visual reasoning abilities of AI. In this work, we perform a preliminary exploration into automated scoring of handwritten student responses to math exams using OpenAI’s GPT-4o model. We evaluate several prompts, analyze failure patterns, and find that GPT-4o’s scoring ability is significantly lower than that of human graders. One recent study also evaluates AI in handwritten math grading [4]. However, they use responses from an optional exam where students may not put in much effort, while we use real final exam responses, yielding a more realistic data source that captures student behavior in actual test-taking settings.
@@ 2 Experimental Setup
Using an IRB-approved process, we collect a dataset of real handwritten final exam responses from a single semester of a probability theory course at a university in the United States. We emailed students from the course asking permission to use their exam responses for research; we use exams from the 18 students who gave consent. We did not collect any demographic information from students. The exam contains 5 questions, each with 3 independent sub-parts, covering the topics of probability estimates, game theory, Markov chains, Bayes nets, and parameter estimation. Each question part also has a rubric that gives credit for partially correct solutions. Students in our sample scored 89.88% on average.
@@ 4 Discussion, Conclusion, and Future Work
In this work, we evaluate the ability of GPT-4o, a powerful multi-modal large vision-language model, to grade real handwritten student responses in college math exams. We find that providing a correct answer and rubric as reference are necessary to improve alignment with human graders, but that GPT-4o still struggles to accurately assign scores for many reasons. In particular, we find that the model struggles to comprehend student solutions, either from i) trouble reading the student’s handwriting, ii) not knowing the true correct solution steps, or iii) incorrectly interpreting the reasoning behind a student’s response. There are many avenues for future work. First, researchers should further assess model weaknesses by prompting the model to transcribe what the student wrote or ask if it feels confident in its score assignment. Second, researchers should investigate if fine-tuning can improve alignment, especially with access to recent open-source vision-language models such as Llama 3.2. Finally, researchers should evaluate visual response grading in other domains, such as handwritten solutions in computer science or even purely visual domains such as visual art or graphic design.

### v2

@@ 1 Introduction
Automated scoring is a key challenge to enable the deployment of open-ended questions at scale. Researchers have studied the problems of automated essay scoring [1] and automated short-answer grading [3] extensively, often using AI. Fine-tuning language models, such as BERT, has been effective on these tasks [6]; others have recently explored prompting large language models (LLMs) for automated scoring [5]. One important setting in automated scoring is student handwriting: in many practice and assessment settings, students write down their solutions to problems on paper, which are then converted to images, common in science, technology, engineering, and math (STEM) fields [2]. This automated scoring task is challenging: compared to scoring textual essays and short answers, images contain rich semantic, visual, and even spatial information on student thought processes, which require significant textual, mathematical, and visual reasoning capabilities from AI. Recent advances in multi-modal foundation models, especially vision-language models, have significantly advanced the textual and visual reasoning abilities of AI. In this work, we perform a preliminary exploration into automated scoring of handwritten student responses to math exams using OpenAI’s GPT-4o model. We evaluate several prompts, analyze failure patterns, and find that GPT-4o’s scoring ability is significantly lower than that of human graders. One recent study also evaluates AI in handwritten math grading [4]. However, they use responses from an optional exam where students may not put in much effort, while we use real final exam responses, yielding a more realistic data source that captures student behavior in actual test-taking settings.
@@ 2 Experimental Setup
Using an IRB-approved process, we collect a dataset of real handwritten final exam responses from a single semester of a probability theory course at a university in the United States. We emailed students from the course asking permission to use their exam responses for research; we use exams from the 18 students who gave consent. We did not collect any demographic information from students. The 120-minute exam contains 5 questions, each with 3 independent sub-parts, covering the topics of probability estimates, game theory, Markov chains, Bayes nets, and parameter estimation. Each question was scored by a single grader using rubrics they wrote to assign credit for partially correct solutions. Student written exam responses contained text, mathematical formulas, and diagrams, all critical to understanding their solutions. Students in our sample scored 89.88% on the exam on average, with many points from partial credit on incorrect answers.
@@ 4 Discussion, Conclusion, and Future Work
In this work, we evaluate the ability of GPT-4o, a powerful multi-modal large vision-language model, to grade real handwritten student responses in college math exams. We find that providing a correct answer and rubric as reference are necessary to improve alignment with human graders, but that GPT-4o still struggles to accurately assign scores for many reasons. In particular, we find that the model struggles to comprehend student solutions, either from i) trouble reading the student’s handwriting, ii) not knowing the true correct solution steps, or iii) incorrectly interpreting the reasoning behind a student’s response. There are many avenues for future work. First, researchers should assess GPT-4o’s performance on sub-tasks, such as transcribing or reasoning over solutions. Second, researchers should investigate if fine-tuning open-source models like Llama 3.2 can improve alignment. Finally, researchers should evaluate visual grading in more domains, such as computer science or visual arts.